# Efficient self-consistent learning of gate set Pauli noise

### v1

@@ VI.2 Learnability of quasi-local noise model
Figure 7: Example of a $3$-qubit quasi-local noise model and the embedding into a complete model. Here, the noise channels for state prep., meas., and all gates are assumed to be $\Omega$-local where $\Omega^{*}=\{\{1,2\},\{2,3\}\}$, with the factor graph shown in the bottom-left side. In other words, $\Lambda^{S}=\Lambda^{S}_{12}\circ\Lambda^{S}_{23}$ where $\Lambda^{S}_{12},~{}\Lambda^{S}_{23}$ are Pauli diagonal maps supporting on $\{1,2\},~{}\{2,3\}$, respectively. Similar for other Pauli noise channels.
@@
  • If the state preparation noise parameters are given by $\{r_{\nu}^{S}:\nu\in\Omega^{S}\}$, where $\Omega^{S}$ is a clique factor set, such that
@@
  • If the noise parameters for some $\mathcal{G}\in\mathfrak{G}$ are given by $\{r_{b}^{\mathcal{G}}:b\sim\Omega^{\mathcal{G}},b\neq I_{n}\}$, where $\Omega^{\mathcal{G}}$ is a clique factor set, such that
@@
Clearly, for a Pauli $P$ with support $s$, the support of $\mathcal{G}(P)$ cannot exceed $\Xi_{\mathcal{G}}(s)$. Lemma 6.5 roughly says that if for all Pauli $P\sim\Omega$, one has $\mathcal{G}(P)\sim\Omega$, then $\Omega$ must be $\mathcal{G}$-covariant. The proof is given in Appendix. D.
@@
Theorem 6(c) confirms that, given certain “goodness” assumptions of the quasi-local noise model, the embedded gauge space is indeed spanned by a subset of SDG’s consistent with the clique factor sets. Similar to the fully local case, one can efficiently compute the reduced gauge space $T_{R}=\mathrm{span}\{\mathcal{Q}^{-1}(\mathfrak{d}_{\nu}):\nu\in\Omega^{S}\}$ by applying the Möbius inverse transformation given in Lemma 6.3.
@@ VI.3 Efficient learning of quasi-local noise model
where we have used the defining property of $\Omega^{M}$ and $\Omega^{\mathcal{G}}$. let us first prove the linear independence of $\mathsf{B}_{R}$. Consider a linear combination of all vectors from ${\mathsf{B}}_{R}$ that yields $\bm{0}$. Since only the second part of ${\mathsf{B}}_{R}$ contains gate noise parameters, and one can show $\{\sum_{b\triangleleft a}\bm{\vartheta}_{b}^{\mathcal{G}}:\mathcal{G}\in\mathfrak{G},~{}a\sim\Omega^{\mathcal{G}}\}$ are linearly-independent (as they can represent $\{\bm{\vartheta}_{a}^{\mathcal{G}}:\mathcal{G}\in\mathfrak{G},a\sim\Omega^{\mathcal{G}}\}$ using the inclusion-exclusion principle, see Appendix D for similar expressions), the coefficients for all vectors from the second part must be zero; On the other hand, because $\{\sum_{\nu\subseteq\mu}\bm{\vartheta}_{\nu}^{M}:\nu\in\Omega^{M}\}$ are also linearly-independent (also thanks to the inclusion-exclusion principle), the coefficients for vectors from the first part must also be zero. This means ${\mathsf{B}}_{R}$ must be linearly independent. On the other hand, Theorem 6(c) implies $\dim(T_{R})=|\Omega^{S}|$. We can see from the expression of ${\mathsf{B}}_{R}$ that $|{\mathsf{B}}_{R}|=\dim(X_{R})-|\Omega^{S}|=\dim(X_{R})-\dim(T_{R})=\dim(L_{R})$. We can thus conclude that ${\mathsf{B}}_{R}$ forms a basis for $L_{R}$. ∎
@@ VII Case study: CZ gates with 1D topology
In this section, we will apply our theory to a concrete gate set where the Control-Z (CZ) gates are the only multi-qubit entangling gates. Such a gate set, augmented with single-qubit gates, are able to conduct universal quantum computation, and is adopted by state-of-the-art experimental platforms (e.g., IBM Quantum [30]555 Note that the Control-NOT (CNOT) and the Echoed Cross-Resonance (ECR) gates are all equivalent to CZ up to single-qubit rotations. Our theory can similarly be applied to them. ). We will discuss the noise learnability and learning algorithms with several different noise ansatz. We hope the results presented here can not only illustrate our theoretical results, but also serve as a new noise characterization protocol for practical use. A quick summary of results is given in Table 2.
@@ VII.3 CZ’s on ring, nearest-neighbor noise
Figure 12: Illustration of the setup in Sec. VII.3 with $n=12$. (a) Qubit topology and noise model. Each circle represents a qubit with their label annotated. Each diamond connects to a maximal clique factor from $\Omega_{*}$, specifying the nearest-neighbor noise model for both SPAM and gate noises. (b), (c) Illustration of $\mathcal{G}_{e},\mathcal{G}_{o}$, respectively.
To begin with, this noise model is not gate-covariant in the sense of Definition 6.4. To see this, consider the Pauli channel $\Lambda^{\prime}(\rho)=\eta X_{2}X_{3}\rho X_{2}X_{3}+(1-\eta)\rho$ with some $\eta<1/2$, which is $\Omega$-local thanks to Lemma D.2. We have $\mathcal{G}_{e}\Lambda^{\prime}\mathcal{G}_{e}^{\dagger}=\eta Z_{1}X_{2}X_{3}Z_{4}\rho Z_{1}X_{2}X_{3}Z_{4}+(1-\eta)\rho$, which is not $\Omega$-local, again by Lemma D.2. Similarly for $\mathcal{G}_{o}$. This means we cannot directly apply Theorem 6(c). In fact, we can show that the reduced gauge space is given by $T_{R}=\mathrm{span}\{\mathcal{Q}^{-1}(\mathfrak{d}_{\{j\}}):j\in[n]\}$, i.e., only the single-qubit depolarizing gauges. Intuitively, any (linear combination of) two-qubit depolarizing gauge transformation will violate the $\Omega$-local noise assumption of either $\mathcal{G}_{e}$ or $\mathcal{G}_{o}$. A formal proof is given in Appendix E.1. Therefore, we obtain the following characterization of the model learnability,
@@
      $\begin{gathered}\{\bm{e}_{a}^{\mathcal{G}_{e}}+\bm{e}_{a^{\prime}}^{\mathcal{G}_{e}}:a\in{\sf P}^{\{2k,2k+1\}},w(a)=2,a^{\prime}=\mathcal{G}_{e}(a)\}\\ \cup\{\bm{e}_{a}^{\mathcal{G}_{o}}+\bm{e}_{a^{\prime}}^{\mathcal{G}_{o}}:a\in{\sf P}^{\{2k+1,2k+2\}},w(a)=2,a^{\prime}=\mathcal{G}_{e}(a)\}\end{gathered}$ (81)
@@
For a proof that the above cycles constitute a cycle basis for $L_{R}^{G}$ and a way to supplement it into a complete basis for $L_{R}$, see Appendix E.2.
@@ VII.4 CZ’s on ring, covariant quasi-local noise
Figure 13: The covariant quasi-local gate noise model from Sec. VII.4. The left and right shows $\Omega^{e}$, $\Omega^{o}$, respectively, with each diamond connecting to a maximal clique factor from $\Omega^{e}_{*}$ (resp. $\Omega^{o}_{*}$).
@@
The factor graphs for each noise model are given in Fig. 13. The fact that this model is gate-covariant follows from Lemma 6.5. Indeed, each maximal clique factor is weight-$4$, containing exactly the support of two neighboring CZ. Consequently, Theorem 6 gives us that
@@
Acknowledgements.
We would like to thank Alireza Seif, Yunchao Liu, Edward Chen, Joseph Emerson, Matthew Graydon, Hsin-Yuan Huang, Ming Yuan, Sisi Zhou, and Yunxiao Zhao for helpful discussions. S.C. & L.J. acknowledge support from the ARO (W911NF-23-1-0077), ARO MURI (W911NF-21-1-0325), AFOSR MURI (FA9550-19-1-0399, FA9550-21-1-0209, FA9550-23-1-0338), NSF (OMA-1936118, ERC-1941583, OMA-2137642, OSI-2326767, CCF-2312755), NTT Research, Packard Foundation (2020-71479).
@@ Appendix A Proof of Theorem 1
The proof follows [15, App. D] with some modification to fit in our framework.
@@ C.2 Proof of converse part of Theorem 5(b)
C.3 Proof of Lemma 5.8
@@
To prove the correctness of the algorithm, we prove the following loop invariant: $(\widetilde{L_{R}^{G}}\cap\mathrm{span}(A))+\mathrm{span}(C)=\widetilde{L_{R}^{G}}$.
@@
Appendix D About quasi-local Pauli channels
@@
The following Lemma explains the equivalence between the quasi-local Pauli channels we consider and the sparse Pauli-Lindblad noise model. We note that basically the same results are given by [14, Theorem SIV.1].
Lemma D.2.
Given an $n$-qubit Pauli channel $\Lambda$ and a clique factor set $\Omega$, if $\Lambda$ can be written in the following concatenating form
@@
Thanks to Lemma D.2, any $\Omega$-local Pauli channel $\Lambda$ can be written in the form of Eq. (111). Conjugate $\Lambda$ with $\mathcal{G}$ yields,
@@
Appendix E Additional proofs for Sec. VII.3
E.1 Gauge in the nearest-neighbor $CZ$ model
Proposition E.1.
@@ E.1 Gauge in the nearest-neighbor $CZ$ model
E.2 Cycle basis for the nearest-neighbor $CZ$ model

### v2

@@ VI.2 Learnability of quasi-local noise model
Figure 4: Example of a $3$-qubit quasi-local noise model and the embedding into a complete model. Here, the noise channels for state prep., meas., and all gates are assumed to be $\Omega$-local where $\Omega^{*}=\{\{1,2\},\{2,3\}\}$. In other words, $\Lambda^{S}=\Lambda^{S}_{12}\circ\Lambda^{S}_{23}$ where $\Lambda^{S}_{12},~{}\Lambda^{S}_{23}$ are Pauli diagonal maps supporting on $\{1,2\},~{}\{2,3\}$, respectively. Similar for other Pauli noise channels.
@@
  • If the state preparation noise parameters are given by $\{r_{\nu}^{S}:\nu\in\Omega^{S}\}$, where $\Omega^{S}$ is a factor set, such that
@@
  • If the noise parameters for some $\mathcal{G}\in\mathfrak{G}$ are given by $\{r_{b}^{\mathcal{G}}:b\sim\Omega^{\mathcal{G}},b\neq I_{n}\}$, where $\Omega^{\mathcal{G}}$ is a factor set, such that
@@
Clearly, for a Pauli $P$ with support $s$, the support of $\mathcal{G}(P)$ cannot exceed $\Xi_{\mathcal{G}}(s)$. Lemma 6.5 roughly says that if for all Pauli $P\sim\Omega$, one has $\mathcal{G}(P)\sim\Omega$, then $\Omega$ must be $\mathcal{G}$-covariant. The proof is given in Appendix. E.
@@
Theorem 11(c) confirms that, given certain “goodness” assumptions of the quasi-local noise model, the embedded gauge space is indeed spanned by a subset of SDG’s consistent with the factor sets. Similar to the fully local case, one can efficiently compute the reduced gauge space $T_{R}=\mathrm{span}\{\mathcal{Q}^{-1}(\mathfrak{d}_{\nu}):\nu\in\Omega^{S}\}$ by applying the Möbius inverse transformation given in Lemma 6.3.
@@ VI.3 Efficient learning of quasi-local noise model
where we have used the defining property of $\Omega^{M}$ and $\Omega^{\mathcal{G}}$. let us first prove the linear independence of $\mathsf{B}_{R}$. Consider a linear combination of all vectors from ${\mathsf{B}}_{R}$ that yields $\bm{0}$. Since only the second part of ${\mathsf{B}}_{R}$ contains gate noise parameters, and one can show $\{\sum_{b\triangleleft a}\bm{\vartheta}_{b}^{\mathcal{G}}:\mathcal{G}\in\mathfrak{G},~{}a\sim\Omega^{\mathcal{G}}\}$ are linearly-independent (as they can represent $\{\bm{\vartheta}_{a}^{\mathcal{G}}:\mathcal{G}\in\mathfrak{G},a\sim\Omega^{\mathcal{G}}\}$ using the inclusion-exclusion principle, see Appendix E for similar expressions), the coefficients for all vectors from the second part must be zero; On the other hand, because $\{\sum_{\nu\subseteq\mu}\bm{\vartheta}_{\nu}^{M}:\nu\in\Omega^{M}\}$ are also linearly-independent (also thanks to the inclusion-exclusion principle), the coefficients for vectors from the first part must also be zero. This means ${\mathsf{B}}_{R}$ must be linearly independent. On the other hand, Theorem 11(c) implies $\dim(T_{R})=|\Omega^{S}|$. We can see from the expression of ${\mathsf{B}}_{R}$ that $|{\mathsf{B}}_{R}|=\dim(X_{R})-|\Omega^{S}|=\dim(X_{R})-\dim(T_{R})=\dim(L_{R})$. We can thus conclude that ${\mathsf{B}}_{R}$ forms a basis for $L_{R}$. ∎
@@ VII Case study: CZ gates with 1D topology
In this section, we will apply our theory to a concrete gate set where the Control-Z (CZ) gates are the only multi-qubit entangling gates. Such a gate set, augmented with single-qubit gates, are able to conduct universal quantum computation, and is adopted by state-of-the-art experimental platforms (e.g., IBM Quantum [32]777 Note that the Control-NOT (CNOT) and the Echoed Cross-Resonance (ECR) gates are all equivalent to CZ up to single-qubit rotations. Our theory can similarly be applied to them. ). We will discuss the noise learnability and learning algorithms with several different noise ansatz. We hope the results presented here can not only illustrate our theoretical results, but also serve as a new noise characterization protocol for practical use. A quick summary of results is given in Table 2.
@@ VII.3 CZ’s on ring, nearest-neighbor noise
Figure 9: Illustration of the setup in Sec. VII.3 with $n=12$. (a) Qubit topology and noise model. Each circle represents a qubit with their label annotated. Each diamond connects to a maximal factor from $\Omega_{*}$, specifying the nearest-neighbor noise model for both SPAM and gate noises. (b), (c) Illustration of $\mathcal{G}_{e},\mathcal{G}_{o}$, respectively.
To begin with, this noise model is not gate-covariant in the sense of Definition 6.4. To see this, consider the Pauli channel $\Lambda^{\prime}(\rho)=\eta X_{2}X_{3}\rho X_{2}X_{3}+(1-\eta)\rho$ with some $\eta<1/2$, which is $\Omega$-local thanks to Lemma E.2. We have $\mathcal{G}_{e}\Lambda^{\prime}\mathcal{G}_{e}^{\dagger}=\eta Z_{1}X_{2}X_{3}Z_{4}\rho Z_{1}X_{2}X_{3}Z_{4}+(1-\eta)\rho$, which is not $\Omega$-local, again by Lemma E.2. Similarly for $\mathcal{G}_{o}$. This means we cannot directly apply Theorem 11(c). In fact, we can show that the reduced gauge space is given by $T_{R}=\mathrm{span}\{\mathcal{Q}^{-1}(\mathfrak{d}_{\{j\}}):j\in[n]\}$, i.e., only the single-qubit depolarizing gauges. Intuitively, any (linear combination of) two-qubit depolarizing gauge transformation will violate the $\Omega$-local noise assumption of either $\mathcal{G}_{e}$ or $\mathcal{G}_{o}$. A formal proof is given in Appendix F.1. Therefore, we obtain the following characterization of the model learnability,
@@
      $\begin{gathered}\{\bm{e}_{a}^{\mathcal{G}_{e}}+\bm{e}_{a^{\prime}}^{\mathcal{G}_{e}}:a\in{\sf P}^{\{2k,2k+1\}},w(a)=2,a^{\prime}=\mathcal{G}_{e}(a)\}\\ \cup\{\bm{e}_{a}^{\mathcal{G}_{o}}+\bm{e}_{a^{\prime}}^{\mathcal{G}_{o}}:a\in{\sf P}^{\{2k+1,2k+2\}},w(a)=2,a^{\prime}=\mathcal{G}_{o}(a)\}\end{gathered}$ (81)
@@
For a proof that the above cycles constitute a cycle basis for $L_{R}^{G}$ and a way to supplement it into a complete basis for $L_{R}$, see Appendix F.2.
@@ VII.4 CZ’s on ring, covariant quasi-local noise
Figure 10: The covariant quasi-local gate noise model from Sec. VII.4. The left and right shows $\Omega^{e}$, $\Omega^{o}$, respectively, with each diamond connecting to a maximal factor from $\Omega^{e}_{*}$ (resp. $\Omega^{o}_{*}$).
@@
The factor graphs for each noise model are given in Fig. 10. The fact that this model is gate-covariant follows from Lemma 6.5. Indeed, each maximal factor is weight-$4$, containing exactly the support of two neighboring CZ. Consequently, Theorem 11 gives us that
@@
Notes added. We became aware of an independent and concurrent work [70] that also studied the learnability of the sparse Pauli-Lindblad noise model and proposed novel learning protocols. We thank Alessio Calzona for communicating with us about their manuscript.
Acknowledgments
We would like to thank Alireza Seif, Yunchao Liu, Edward Chen, Joseph Emerson, Matthew Graydon, Hsin-Yuan Huang, Ming Yuan, Sisi Zhou, and Yunxiao Zhao for helpful discussions. S.C. & L.J. acknowledge support from the ARO(W911NF-23-1-0077), ARO MURI (W911NF-21-1-0325), AFOSR MURI (FA9550-21-1-0209, FA9550-23-1-0338), DARPA (HR0011-24-9-0359, HR0011-24-9-0361), NSF (ERC-1941583, OMA-2137642, OSI-2326767, CCF-2312755, OSI-2426975), and Packard Foundation (2020-71479).
@@ Appendix A Proof of Theorem 6
The proof follows [16, App. D] with some modification to fit in our framework.
@@ C.2 Proof of converse part of Theorem 10(b)
Appendix D Relative-precision learning for 2-qubit gates with fully-local noise
We consider the second type of algorithms that learn gate noise parameters to relative precision. The key is to construct a reduced cycle basis for the gate noise parameter space $L_{R}^{G}=\mathcal{Q}^{\sf T}(L\cap X^{G})$. Note that, one might think
We focus on the case that every $\mathcal{G}\in\mathfrak{G}$ is supported non-trivially on at most $2$ qubits. Since many existing quantum computing platform uses two-qubit entangling gates such as CZ as a primitive, we expect our method to be practically useful.
Algorithm 3  [Finding a basis for $\widetilde{L_{R}^{G}}$].
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Repeat step 2 and 3 until no $\bm{c}$ can be found, return $C$.
Lemma D.1.
Algorithm 3 outputs a basis for $\widetilde{L_{R}^{G}}$.
This is a standard linear algebraic procedure. For completeness, a proof is given below.
@@ Appendix D Relative-precision learning for 2-qubit gates with fully-local noise
We prove the following loop invariant: $(\widetilde{L_{R}^{G}}\cap\mathrm{span}(A))+\mathrm{span}(C)=\widetilde{L_{R}^{G}}$.
@@
Definition D.2.
For $\bm{r}\in X_{R}^{G}$, $\partial(\bm{r})$ is called its boundary.
Lemma D.3.
Proof.
$\Leftrightarrow$ $\quad\partial(\bm{r})=\bm{0}.$
In the first stage, we look for $\bm{c}=\sum_{a\in\sf P^{2}\backslash\{I_{2}\},\mathcal{G}\in\mathfrak{G}}r_{a}^{\mathcal{G}}\bm{\vartheta}_{a}^{\mathcal{G}}$ with the property that all non-zero $r_{a}^{\mathcal{G}}$ share the same $\mathrm{supp}(\mathcal{G})$. That is, we group the base vectors $\bm{\vartheta}_{a}^{\mathcal{G}}$ by $\mathrm{supp}(\mathcal{G})$ and consider them separately. Fix a common support $s$, the task simplifies substantially since Lemma D.3 simplifies into a requirement on $2$ dimensions (the common support, which is of size $2$). Then we may perform the analysis on the pattern transfer subgraph (the union of the pattern transfer subgraphs of $\hat{\mathcal{G}}$ for all $\mathcal{G}\in\mathfrak{G}$ with $\mathrm{supp}(\mathcal{G})=s$). In the following we consider the possible shapes of the pattern transfer subgraph. Since self loops can be indentified and picked as reduced cycle basis easily at this stage, we do not need to explicitly consider them.
Lemma D.4.
Proof.
From Figure 11 one can easily see that the union of pattern transfer subgraphs still has only $3$ possible shapes. The important message of Lemma D.4 is that for any edge in the pattern transfer subgraph, there is another edge of the reverse direction.
$\Downarrow$
(a) There exists two type-1 boundaries with the same support.
{NiceTabular}
$\Downarrow$
[…]
(b) Chasing procedure at both end terminates with a type-1 boundary.
{NiceTabular}
$\Downarrow$
[…]
(c) A chasing procedure ends with a loop.
Now the algorithm proceeds to the second stage, where we find reduced cycles across different supports. This stage is illustrated in Figure 12. According to Lemma D.3 we associate the base vectors in $A$ with their boundaries. Thanks to Lemma D.4, there are only two types of boundaries: Type 1 boundary takes ±$\pm 1$ at one entry and 00 at other places, corresponding to edges (10,11) or (01,11) in Fig. 11. Type $2$ boundary takes $1$ at one entry, $-1$ at another entry, and 00 at other places, corresponding to edge (10,01) in Fig. 11. Moreover, if we consider the edge of reverse direction, the corresponding boundary is negated. In the following we will find some candidate $\bm{c}=\sum\mathrm{sgn}_{i}\bm{\vartheta}_{i}$ and then prove that it is a reduced cycle when all the signs are positive. When some signs are negative, we revert some edges to make the signs positive. This is actually a simplification and hides the following fact: though $\bm{c}$ itself is not a reduced cycle, $\tilde{\bm{c}}=\bm{c}+\sum_{\mathrm{sgn}_{i}=-1}(\bm{\vartheta}_{i}+\bm{\vartheta}_{i}^{\prime})$ ($\bm{c}$ with edge directions reverted) is a reduced cycle. Moreover, since the $\bm{\vartheta}_{i}+\bm{\vartheta}_{i}^{\prime}$s are already in $C$, putting $\bm{c}$ or $\tilde{\bm{c}}$ into $C$ is equivalent in terms of making $C$ a basis for $\widetilde{L_{R}^{G}}$.
According to the proof of Lemma D.1, if there are remaining degree of freedom in $\widetilde{L_{R}^{G}}$, there should exist a non-zero vector $\bm{c}^{\prime}\in\widetilde{L_{R}^{G}}\cap\mathrm{supp}(A)$. Then $\partial(\bm{c}^{\prime})=\bm{0}$. We can find such $\bm{c}^{\prime}=\sum_{i}r_{a_{i}}^{\mathcal{G}_{i}}\bm{\vartheta}_{a_{i}}^{\mathcal{G}_{i}}$ (the summation is only over non-zero terms) via simple linear algebra on the boundaries of the edges in $A$. The $r_{a_{i}}^{\mathcal{G}_{i}}$ found is rational and thus can be made integral by multiplying them by the least common multiple of the denominators.
Appendix E About quasi-local Pauli channels
@@
The following Lemma explains the equivalence between the quasi-local Pauli channels we consider and the sparse Pauli-Lindblad noise model. We note that basically the same results are given by [15, Theorem SIV.1].
Lemma E.2.
Given an $n$-qubit Pauli channel $\Lambda$ and a factor set $\Omega$, if $\Lambda$ can be written in the following concatenating form
@@
Thanks to Lemma E.2, any $\Omega$-local Pauli channel $\Lambda$ can be written in the form of Eq. (115). Conjugate $\Lambda$ with $\mathcal{G}$ yields,
@@
Appendix F Additional proofs for Sec. VII.3
F.1 Gauge in the nearest-neighbor $CZ$ model
Proposition F.1.
@@ F.1 Gauge in the nearest-neighbor $CZ$ model
F.2 Cycle basis for the nearest-neighbor $CZ$ model